• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Meral

    Log entry time 08:01:53 on August 17, 2006

    Entry number 180523

    keyword=OWL Shift/Run Summary

    Shift Summary
    00:00- Continuing Kin Point 25
    - Current Setting LD2, 5%,
    00:09- VESDA Hall Wall South Alarm
    00:15- JT valve settings changed
    00:26- CODA died, reset the crate
    01:25- FSD tripped frequently
    01:48- no beam, transport problems
    02:27- beam is back, bmp fluctuating
    03:04- MCC called optics guy is here
    05:22- beam is back, but fluctuating
    05:45- MCC called saying they would be taking beam away, but did not
    follow through
    6:20- MCC called to say we have production beam again
    7:20- MCC called sending tuned beam for a few minutes



    Run Summary
    Run Num, Num Events, Tot Current
    2838/21811 1.9M, 13.3M 10uA
    2839/21182 left coda died, junk
    2840/21183 1.9M, 15.2M, 10uA
    2841/21184 1.9M, 17.1M 10uA
    2842/21185 unstable beam!
    2843/21186 1.9M, 19M 10uA
    2844/21187 unstable beam!
    2845/21188 1.9M, 20.9M 10uA ? not production beam ?
    2846/21189 1.9M, 22.8M 10uA
    2847/21190 1.9M, 24.7M 10uA
    2848/21191 1.7M, 26.4M 10uA , stopped for tuned beam